The profit from the sale of x units is P(x)=80x−1200−x^2

a) How many units do they need to sell in order to break even? Enter your answers separated by a comma.

   

b) What production level maximizes profit, and what is the profit?


Production level =  

Profit = $
Answer #1

2 P(x) = 80% -1200 - To achieve Break Even points, profit= o te Tofal Revenue = Total costs so Pla=0 → 807-1200-720 102 - 80x30 production level = 40 know 40 is the maximizes profits achieve , u TO we Gel from equi dp(x) = 80-21 da 7 d?p(x) = -2 a az